OpenStudy (anonymous):

Find the volume of a cylinder with a diameter of 10 inches and a height that is three times the radius. Use 3.14 for pi and round your answer to the nearest tenth. (Hint: You may only enter numerals, decimal points, and negative signs in the answer blank)

OpenStudy (anonymous):

@Haseeb96 @Isaiah.Feynman @shamil98

OpenStudy (isaiah.feynman):

@prowrestler can you find the radius of the cylinder?

OpenStudy (shamil98):

Use the formula : \[volume = {\pi}r^2h\]

OpenStudy (shamil98):

The radius is half of the diameter, btw.

OpenStudy (anonymous):

@Isaiah.Feynman I think I can

OpenStudy (isaiah.feynman):

Find it.

OpenStudy (haseeb96):

d=2r so r=5inches v=pi r^2 h here h is three time the radius so , h =3r h=3(5) h=15 now put it in the formula of volume of cylinder

OpenStudy (anonymous):

@Isaiah.Feynman the radius would be 5

OpenStudy (shamil98):

yes the radius is 5

OpenStudy (shamil98):

The height is three times the radius so what is it?

OpenStudy (anonymous):

@Haseeb96 I got 1178.097245

OpenStudy (haseeb96):

v=3.14 x 25 x15 V=1177.5inches^3

OpenStudy (anonymous):

@Haseeb96 is that the answer: 1177.5inches^3

OpenStudy (haseeb96):

yeah u got the correct answer

OpenStudy (shamil98):

\[\huge v = 3.14(25)(15)\]

OpenStudy (isaiah.feynman):

I THINK THIS QUESTION IS TOO ELEMENTARY FOR YOU ALL TO SCRAMBLE TO SHOW YOUR SOLUTIONS.

OpenStudy (anonymous):

wait so is it 1177.5inches

OpenStudy (shamil98):

yes

OpenStudy (anonymous):

thank you all

OpenStudy (anonymous):

ok I have some more

OpenStudy (haseeb96):

how much are left?

OpenStudy (anonymous):

5

OpenStudy (haseeb96):

okay cool